Evaluate sinē 172° - 15cosē 32° to four significant digits.
Answer by addingup(3677)

Rewrite:
[(sin172)^2]-[15(cos32)^2]
= 0.1392^2-15(0.8480)^2
= 0.0193-15(0.7191)
= 0.0193-10.7865
= -10.7672 where the last 2 decimals may change slightly because of rounding.
:
John